Question 3 [1.3 pts] What type of proton NMR signal will proton Hc give rise to? Hc Ho Question 4 [1.3 pts] What type of proton NMR signal will proton Ho give rise to?
Answer #1


according to the (N+1) rule where N = neighbouring protons

in Hc it is having the n = 6 so 6+1 = 7 septet

in Hd it is having the n = 4 so 4+1 = 5 pentet or quintet